# Exactly solvable model of the square-root price impact dynamics under the long-range market-order correlation

### v2

@@ Appendix A Appendix
In this Appendix, the outline of our theoretical calculations is given. Here we simplify the notation based on Ref. [19], by rewriting a single-trader contribution $\Delta m^{(j)}(t)$ as the displacement of a Levy-walk particle $x(t)$.
A.1 Appendix A: the Lévy-walk theory
@@
$\left\langle e^{-{\rm i}kx_{\rm tot}}\right\rangle=\prod_{i=1}^{M}\left\langle e^{-{\rm i}kx^{(i)}}\right\rangle=\left[\left\langle e^{-{\rm i}kx^{(i)}}\right\rangle\right]^{M}$ (15)
because $x^{(i)}$ is independent of $x^{(j)}$ for $i\neq j$. Thus, studying a single-trader market impact is sufficient to understand the statistics of the total market impact. For example, we obtain $\langle x^{2}_{\rm tot}\rangle=M\langle(x^{(i)})^{2}\rangle$ because $\langle x^{(i)}\rangle=0$.
A.2 Appendix B: the mean-squared displacement
@@
A.3 Appendix C: the inverse-cubic law based on the large-deviation principle
@@
with some coefficients $A^{\prime}$ and $B^{\prime}$. By using the Tauberian theorem [19, 29], we obtain the power-law tail (8).
A.4 Appendix D: the inverse-cubic law based on a long resting-time approximation
@@
A.5 Appendix E: the inverse-cubic law based on heuristic arguments

### v3

@@ I Methods
In this section, the outline of our theoretical calculations is given. Here we simplify the notation based on Ref. [15], by rewriting a single-trader contribution $\Delta m^{(j)}(t)$ as the displacement of a Levy-walk particle $x(t)$. More detailed calculations are presented in Supplementary information.
I.1 Appendix A: the Lévy-walk theory
@@
$E\left[e^{-{\rm i}kx_{\rm tot}}\right]=\prod_{i=1}^{M}E\left[e^{-{\rm i}kx^{(i)}}\right]=\left\{E\left[e^{-{\rm i}kx^{(i)}}\right]\right\}^{M}$ (12)
because $x^{(i)}$ is independent of $x^{(j)}$ for $i\neq j$. Thus, studying a single-trader market impact is sufficient to understand the statistics of the total market impact. For example, we obtain $E[x^{2}_{\rm tot}]=M\times E[(x^{(i)})^{2}]$ because $E[x^{(i)}]=0$.
I.2 Appendix B: the mean-squared displacement
@@
I.3 Appendix C: the inverse-cubic law
Figure 4: (a) Complementary cumulative distribution function for the price changes, showing the inverse-cubic law $P_{>}(\Delta m)\propto(\Delta m)^{-\beta}$ with $\beta=3$. We assumed the following parameters: $\delta=1/2$, $\alpha=1.5$, $\tau_{r}=10^{5}$, $t=10^{4}$, and $M=1$. See the inset for the corresponding semi-log plot. (b) Volatility ACF in our model, numerically showing the long memory $C_{V}(\tau)\propto\tau^{-\zeta}$ with $\zeta\in(0,1)$ for the parameters $M=1$, $\delta=0.5$, $\tau_{r}=10$, and $t=1000$. Empirically, the power-law exponent obeys $\zeta\approx\alpha-1$. More detailed numerical investigation is presented in the SI.
Our model exhibits the inverse-cubic law (7a), or equivalently,
$P_{>}(\Delta m):=\int_{\Delta m}^{\infty}P(|\Delta m^{\prime}|){\rm d}\Delta m^{\prime}\propto(\Delta m)^{-\beta},\>\>\>\beta:=\frac{\alpha}{\delta},$ (14)
where $P_{>}(\Delta m)$ is the complementary cumulative distribution function for the price changes $\Delta m:=m(t)-m(0)$. See Fig. 4(a) for the numerical simulation. Let us derive this formula by three methods below.
I.3.1 Derivation based on the large deviation principle
@@
with some coefficients $A^{\prime}$ and $B^{\prime}$. By using the Tauberian theorem [15, 26], we obtain the power-law tail (7a).
I.3.2 Derivation based on a long resting-time approximation
@@
I.3.3 Heuristic derivation
@@
II Appendix D: Volatility clustering
We also numerically study the long memory of the volatility in our model. The model parameter was given by $M=1$, $\delta=1/2$, $\tau_{r}=10$, and $t=1000$. In Fig. 4(b), we plot the volatility ACF
$C_{V}:=\frac{E[(\sigma^{2}(0,t)-\mu_{\sigma^{2}_{t}})(\sigma^{2}(\tau,\tau+t)-\mu_{\sigma^{2}_{t}})]}{E[(\sigma^{2}(0,t)-\mu_{\sigma^{2}_{t}})^{2}]},$ (22)
with $\sigma^{2}(t,t+\tau):=\{p(t+\tau)-p(t)\}^{2}$ and $\mu_{\sigma^{2}_{t}}:=E[\sigma^{2}(0,t)]$. The figure shows the power-law decay $C_{V}(\tau)\propto\tau^{-\zeta}$ with $\zeta\in(0,1)$. We numerically find the relationship
$\zeta\approx\alpha-1.$ (23)
We leave the mathematical derivation of this empirical formula as a future issue.
Code availability
Simulation codes are available as our SI.
Acknowledgements
Acknowledgements.
We thank for the fruitful comments by J.-P. Bouchaud and M. Guillaume. YS was supported by JSPS KAKENHI (Grant No. 24KJ1328). KK was supported by JSPS KAKENHI (Grant Nos. 21H01560 and 22H01141) and the JSPS Core-to-Core Program (Grant No. JPJSCCA20200001).
Author contribution
YS developed the numerical program code and contributed to the analytical calculations for the mean-squared displacement. KK designed and managed the project, and contributed the analytical calculations, particularly for the inverse-cubic law. Both YS and KK contributed to writing the manuscript and approved its final content.
Conflict of interest
We declare no financial conflict of interest.